Hello,
Please help me with this: 4/31=5/x
Thank you

Answer by stanbon(75887)

4/31=5/x
Invert both sides to get:
x/5 = 31/4
Multiply both sides by 5 to get:
x=5(31/4) = 155/4
Cheers,
Stan H.